OpenStudy (anonymous):

write an equation in slope-intercept form of a line containing the point (9,-7) with a slope equal to 3. a)y=3x+20 b)y=3x+34 c)y=3x-20 d)y=3x-34

OpenStudy (anonymous):

\[y+7=3(x-9)\] multiply using dist law \[y+7=3x-27\] subtract 7 \[y=3x-34\]